A circular loop in the plane of the paper lies in a 0.68 T magnetic field pointing into the paper.

What is the magnitude of the average induced emf?

If the coil resistance is 2.1 Ω , what is the average induced current?
Answer #1

We know that the emf is induced when magnetic flux changes across the field
that is
E = -d(Phi)/dt
where phi is magnetic flux phi = BA Cos theta
since the angle is zero between area vector and magnetic field therefore
Cos theta = 1
anf magnetic flux Phi = BA
if diameter changes from d1 to d2 in time t
then area change from
A1 = (Pi/4)(d1)2  to A2 = (Pi/4)(d2)2
Therefore
Induced emf = B.(A1 - A2)/t
where t is the time to change the diameter.
E =  B.(A1 - A2)/t
We know that the
Current = induced emf / resistance = E / R = E /2.1 ampere
Please note that the information provided by you is not complete, so please put down the values as given in the question.